# Transverse beam instabilities in low-emittance booster synchrotrons

### v1

@@ I Introduction
However, this may no longer be acceptable for the new generation boosters built for 4th Generation Light Sources (4GLSs) since they need to provide a much lower emittance beam at extraction to match the usually small dynamic aperture of 4GLS storage rings. For SOLEIL II [1, 2, 3], the upgrade project towards a 4GLS of SOLEIL, for example, the horizontal beam emittance at the new booster’s extraction will be , compared to in the existing booster. This new value is almost identical to the natural emittance in the present 3GLS storage ring of SOLEIL. Such a low emittance value of the new booster imposes that the beam cannot tolerate instabilities that will explode the emittance as much as it could have for 3GLSs. Thorough collective effects studies in a 4GLS booster, therefore, need to be carried out.
@@
This article puts together studies on TSBI and TCBI in a low-emittance booster synchrotron that will serve as an injector for a 4GLS storage ring. We present an observation of a sawtooth instability in the single-bunch regime during the booster ramp process, as well as a demonstration that longitudinal synchrotron damping can also give a damping effect in the transverse planes. We also prove that the increase of instability threshold under the presence of nonlinearities is due to the Landau damping effect, which can be effective against both TSBI and TCBI. Lastly, we show that broad-band impedance can result in a suppression of TCBI at low chromaticity. While the effect of these elements might be well known individually in storage rings, we present a comprehensive study in the context of a booster machine. This work demonstrates the relevance of these effects to the design of future ultra-low emittance booster synchrotrons.
The SOLEIL II booster is used as an example for this study. It has a two-fold symmetry 16-Bend Achromat (16BA) lattice [7, 8, 9] designed to fit in the existing booster tunnel. Its main parameters can be found in Table 1. Note that the emittance, bunch length, and energy spread in this table are the equilibrium values at the corresponding energy. As the beam can never reach the equilibrium at 150 MeV during the ramp, the values appear in the 150 MeV column in this table, therefore, do not show the beam parameters at injection. Instead, the latter is listed in Table 2. Two different bunch lengths at injection are used depending on the regime being studied. This is in accordance with the two operation modes of the upstream linear accelerator: the short-pulse mode for single-bunch, and the long-pulse mode for multibunch operation mode of the storage ring [10]. The injected beam also has an equal emittance in the horizontal and vertical planes. The booster’s impedance model can be consulted in [10]. This study can be of great relevance to other low-emittance boosters, especially those foreseen with a high-charge operation mode for a swap-out injection scheme or with a small vacuum chamber radius since these configurations can give an immense rise to the machine’s coupling impedance, hence possibly strong collective effects.
@@
This article is separated into two main parts: Section II presents the study in the transverse single-bunch regime, and Section III presents the multibunch one. In the first part, we show the use of a stability diagram to assess the beam’s stability at injection, followed by simulations with an energy ramp that can reveal emittance growth along the ramp, if it exists. We also present in this part the observation of sawtooth instability and the effect of longitudinal damping on TSBI. In the second part, a general trend of the TCBI along the ramp in a booster is given via a simple mathematical model. Then, tracking simulation results are presented and prove a certain level of accuracy of this model. In this part, we will see that TCBI is mitigated not only by ADTS but also by broad-band impedance, where the latter is generally unwanted because it can excite single-bunch instability.
@@
Rms H&V emittance | 0.68 (round beam) | ||
@@ II.1 At injection energy
One of the most general mitigation mechanisms for most collective instabilities is synchrotron radiation, but it is very weak at the energy scale of hundreds of MeV, as we will see in Section III.1. However, a low-emittance booster will likely have strong sextupole magnets to correct the highly negative chromaticity due to strong focusing magnets. The strong nonlinear fields of the sextupoles will lead to a highly nonlinear tune shift, i.e., strong ADTS. As a consequence, the beam will have a large incoherent tune spread (rms tune spread ), which means that the coherent motion of the particles inside the beam is broken, leading to a weakening of collective effects. The result can be that the instability totally vanishes or its magnitude is attenuated. This mechanism is known as Landau damping [14, 15] and it has always been a common mitigation method for hadron machines [16, 17, 18] since synchrotron radiation is usually too weak for such particles.
@@
We will, therefore, compare instability growth rates to the limit of the Landau damping to assess the stability at the injection energy. To do so, one can draw a stability diagram which maps between the complex tune shift of the beam when ADTS is present and that of the case where ADTS is absent. The diagram is drawn on a space between the imaginary part of the complex coherent tune shift and its real part . Physically interpreted, the former represents an instability growth rate, while the latter represents the coherent tune shift, the Fourier transform of the center-of-mass position of the bunch. The curve on the diagram shows the solution to the dispersion relation [19] where in the case of a bunched beam with ADTS. This stability boundary should agree with the tracking simulation results; see, e.g., [20].
To determine the beam’s stability, we must track the beam without ADTS to observe the intrinsic properties without the intervention of Landau damping, then extract the coherent tune shift and the instability growth rate. If the data point falls in the area under the stability limit curve, it means that the instability can be stabilized since the growth rate is less than the Landau damping rate. Otherwise, if the data point falls outside this area, the beam will be unstable.
@@
Figure 1 shows the stability diagram of the SOLEIL II booster at 150 MeV, overlayed with the transverse complex tune shift data extracted from tracking simulations with different bunch charges. The horizontal axis is the mode tune shift where is the coherent tune obtained from the simulation, is the zero current coherent tune in the vertical plane, is the head-tail mode number, and is the synchrotron tune. This is a standard way to visualize the data, instead of showing the raw coherent tune shift, it shows the detuning from the corresponding head-tail mode in the unit of . In this plot, the charge is varied between . The data point for each chromaticity 111The chromaticities throughout this article are non-normalized chromaticities. evolves from the lower center of the diagram at a low charge, then climbs up vertically as the charge increases until it reaches the stability limit. It finishes in the upper left corner area at high charge at last. We can see that, at the nominal charge of 0.5 nC, the beam finds itself lying outside the stability limit with a growth rate in the order of while the maximum Landau damping rate is about .
@@
According to this result, the beam’s stability cannot be assumed for the rest of the ramp. However, as the head-tail growth rate decreases with energy, while the synchrotron radiation damping rate increases, it is possible that the instability will not appear at all. Therefore, simulations with the energy ramp are needed to evaluate the instability along the ramp.
@@ II.2 Along the ramp
We perform tracking simulations with mbtrack2, which uses the one-turn maps tracking method. A variation of beam energy is added to predict the realistic evolution of the beam throughout the ramp. The principle is to update, at each turn, the beam energy and equilibrium parameters such as energy loss per turn, radiation damping time, natural energy spread, natural emittance, etc. The beam energy variation itself is calculated via the energy gain per turn equation in relation to the dipole field ramping rate, and the equilibrium parameters are calculated using their proportionality to the beam energy. More details on the booster ramp model can be consulted in [21, 10].
@@
For the nominal charge in Fig. 2(a), it turns out that the beam is well stable and no sign of instability can be observed regardless of chromaticity. The vertical emittance is damped all the way from injection and reaches the designed value 0.5 nm rad at the extraction, which is the desirable situation. This result is rather surprising since some emittance increase could have been expected due to the high growth rate above the Landau stability limit shown in Fig. 1. The explanation for this is that even though Landau damping cannot totally suppress the instability, the growth rates are significantly reduced by the presence of ADTS, in addition to the reduction due to the increasing beam energy. This can be seen in Fig. 3 where growth rates extracted from the tracking simulations with and without ADTS are compared. For a certain chromaticity, the ADTS can greatly diminish the instability growth rate even at zero chromaticity where the Transverse Mode-Coupling Instability (TMCI), which is extremely strong once emerged, takes place. At 0.5 nC for example, the growth rates drop from , as also seen in Fig. 1, to which is less than or in the same order as the booster’s half-cycle rate . This means that the instability will never have enough time to build up before reaching the extraction, and the beam is simply damped by synchrotron radiation as the energy increases.
The growth rate reduction outside the stability limit can actually be viewed as a distortion of the complex tune shift space due to the presence of nonlinearity as we can see in Fig. 4. Here, the same stability limit shown previously is drawn in blue. But, this is only a curve that corresponds to Im[]=0, the definition of a stability boundary. It means that the growth rate along this curve is zero, and we can also call it an isoline of the zero growth rate. There are, in fact, an infinite number of isolines that show other growth rates (see e.g [22]). Clearly seen in the plot, the neighboring space of the stability limit is distorted. If we have a complex tune shift falling, for example, at the point marked as “x” in Fig. 4, it means that the growth rate of without ADTS will be reduced to about under the ADTS.
@@ II.2.1 Sawtooth instability
Unique to the case of maximum charge, an unexpected sawtooth feature was observed at chromaticities 3 to 5. Further investigations were, therefore, carried out by simulating TSBI at different energies and the growth rates were compared to the analytical synchrotron radiation damping rate. It should be noted that, in this case, synchrotron radiation was not included in the transverse planes during the tracking in order to speed up the calculation and because we were interested in the pure instability growth rate. It was, however, included in the longitudinal plane to retain the particles within the rf bucket.
The investigation results are shown in Fig. 5. Here, the growth rate of a 1.5 nC bunch at different chromaticities is determined at various points along the ramp discretely, meaning that each point represents an independent simulation with a fixed beam energy. We can see that the growth rates of the high chromaticity group (3 to 5) always remain above the radiation damping rate, while those of the low chromaticity group (0 to 2) drop below the latter at some point. Note that the growth rate at zero chromaticity has dropped below since the first 20 ms. As we have seen in Fig. 2(b) that only in the high chromaticity group did the sawtooth behavior occur, it is clear that the growth rate being stronger than radiation damping is the reason for the regrowing emittance. Another factor that supports this regrowth is the ineffectiveness of the ADTS at this point of the ramp due to the small emittance. However, as the emittance increases from the instability, the effect of ADTS becomes stronger and can damp the beam again. This mechanism repeats itself as long as the growth rate is larger than the radiation damping rate and it is the explanation to the sawtooth-like shape of the emittance. Such a behavior has also been observed both in simulations and measurements carried out at MAX IV [23], when the influence of ADTS on TMCI was being studied.
@@ II.2.2 Effect from the longitudinal damping
The results in Fig. 6 agree with the study in [25] where the effect of longitudinal damping on TCBI is investigated. In this reference, it is also shown that this effect is stronger when the longitudinal damping rate is larger. However, in our case, when we look at different chromaticities at the nominal longitudinal damping rate, the results do not follow the same conclusion as in this reference. We found that chromaticity 5 is the least affected by the longitudinal damping, and the lower the chromaticity, the stronger the effect. Meanwhile, in the reference, the effect of longitudinal damping is more important at higher chromaticity.
A benchmark on the growth rate as a function of chromaticity was done with a Vlasov solver DELPHI [26] and the result is shown in Fig. 7. Both codes were given with the same parameters and the same bunch length at each energy. The only difference is that DELPHI does not have radiation damping in any plane, while mbtrack2 still has one in the longitudinal plane. It shows clearly that the longitudinal radiation damping plays a role only up to chromaticity 2.
@@ III.1 General trend from analytical estimation
where is the average optic beta function, is the angular revolution frequency, is the total beam current, is the beam energy in eV, is the machine radius, is the chamber radius, is the speed of light, is the impedance of free space, is the fractional part of the betatron tune, and is the conductivity of the chamber wall. Apart from the vacuum chamber radius dependence, we can also see that the RWI growth rate is inversely proportional to the beam energy, .
@@
is also shown. Note that, here, we are interested in the zero chromaticity circumstance, which can only provoke the mode . Theoretically, the longitudinal damping effect can then be neglected. A positive total growth rate means that instability is stronger than synchrotron radiation, and the beam’s center-of-mass (CM) oscillation can grow exponentially in the transverse plane, which is the sign of RWI. The point where the total growth rate goes into the negative side, therefore, marks the point where synchrotron radiation is strong enough to overtake RWI and possibly stabilize the beam. It can be seen clearly in this figure that synchrotron radiation is ineffective during the first half of the ramp, resulting in possible instability.
To quickly predict the beam’s behavior, we define a simple equation to describe the beam’s CM transverse oscillation amplitude as a function of time :
@@
This result gives a general trend of RWI in every booster at zero chromaticity when there are no other stabilizing mechanisms involved, such as Landau damping, head-tail damping, or a transverse feedback system. The curve’s height depends on each machine’s ramp speed: the faster the ramp, the sooner synchrotron radiation can suppress RWI, and the smaller the peak oscillation will be. Another important factor is the vacuum chamber radius. For 3GLS boosters whose chamber radius is typically large, the RWI growth rate is small, so the CM oscillation is relatively small as a consequence.

### v2

@@ I Introduction
However, this may no longer be acceptable for the new generation boosters built for 4th Generation Light Sources (4GLSs) since they need to provide a much lower emittance beam at extraction to match the usually small dynamic aperture of 4GLS storage rings. Taking for example SOLEIL II [1, 2, 3], the upgrade project towards a 4GLS of SOLEIL, the horizontal beam emittance at the new booster’s extraction will be , compared to in the existing booster. This new value is almost identical to the natural emittance in the present 3GLS storage ring of SOLEIL. Such a low emittance value of the new booster imposes that the beam cannot tolerate instabilities that will explode the emittance as much as it could have for 3GLSs. Thorough collective effects studies in a 4GLS booster, therefore, need to be carried out.
@@
This article puts together studies on TSBI and TCBI in a low-emittance booster synchrotron that will serve as an injector for a 4GLS storage ring. We present an observation of a sawtooth instability in the single-bunch regime during the booster ramp process, as well as a demonstration that longitudinal synchrotron radiation damping can also give a damping effect in the transverse planes. We also prove that the increase of instability threshold under the presence of nonlinearities is due to the Landau damping effect, which can be effective against both TSBI and TCBI. Lastly, we show that broad-band impedance can result in a suppression of TCBI at low chromaticity. While the effect of these elements might be well known individually in storage rings, we present a comprehensive study in the context of a booster machine. This work demonstrates the relevance of these effects to the design of future ultra-low emittance booster synchrotrons.
The SOLEIL II booster is used as an example for this study. It has a two-fold symmetry 16-Bend Achromat (16BA) lattice [7, 8, 9] designed to fit in the existing booster tunnel. Its main parameters can be found in Table 1. Note that the emittance, bunch length, and energy spread in this table are the equilibrium values at the corresponding energy. As the beam can never reach the equilibrium at 150 MeV during the ramp, the values appear in the 150 MeV column in this table, therefore, do not show the beam parameters at injection. These are instead shown separately in Table 2 for clarity. Two different bunch lengths at injection are used depending on the regime being studied. This is in accordance with the two operation modes of the upstream linear accelerator: the short-pulse mode for single-bunch, and the long-pulse mode for multibunch operation mode of the storage ring [10]. The injected beam also has an equal emittance in the horizontal and vertical planes. The booster impedance model can be consulted in [10]. This study can be of great relevance to other low-emittance boosters, especially those foreseen with a high-charge operation mode for a swap-out injection scheme or with a small vacuum chamber radius since these configurations can give an immense rise to the machine coupling impedance, hence possibly strong collective effects.
@@
This article is separated into two main parts: Section II presents the study in the transverse single-bunch regime, and Section III presents the multibunch one. In the first part, we show the use of a stability diagram to assess the beam stability at injection, followed by simulations with an energy ramp that can reveal emittance growth along the ramp, if it exists. We also present in this part the observation of sawtooth instability and the effect of longitudinal damping on TSBI. In the second part, a general trend of the TCBI along the ramp in a booster is given via a simple mathematical model. Then, tracking simulation results are presented and prove a certain level of accuracy of this model. In this part, we will see that TCBI is mitigated not only by ADTS but also by broad-band impedance, where the latter is generally unwanted because it can excite single-bunch instability.
@@
| 4 rms H&V emittance | 0.68 (round beam) | ||
@@ II.1 At injection energy
One of the most general mitigation mechanisms for most collective instabilities is synchrotron radiation, but it is very weak at the energy scale of hundreds of MeV, as we will see in Section III.1. However, as a low-emittance booster will likely have strong sextupole magnets, their strong nonlinear fields will lead to a highly nonlinear tune shift, i.e., strong ADTS. The beam will then have a large incoherent tune spread (rms tune spread ) as a consequence, leading to suppression or weakening of beam instabilities via Landau damping. The result can be that the instability totally vanishes or its magnitude is attenuated. This mechanism is known as Landau damping [14, 15, 16] and it has always been a common mitigation method for hadron machines [17, 18, 19] since synchrotron radiation is usually too weak for such particles.
@@
We will, therefore, compare instability growth rates to the limit of the Landau damping to assess the stability at the injection energy. To do so, one can draw a stability diagram which maps between the complex tune shift when ADTS is present and that in the case where ADTS is absent. In this article, we will denote the first quantity as and the second as to highlight the fact that the latter is obtained in the absence of ADTS, i.e., without the nonlinearity. To construct a stability diagram, we need to extract from simulations and plot the imaginary part against their real part. Physically interpreted, the former represents the instability growth rate, while the latter represents the coherent tune shift, the Fourier transform of the center-of-mass position of the bunch. Then, on the same space, we overlay a curve called the stability limit which marks where zero growth rate will be if ADTS is present. If the data point falls under the curve, it means that the beam will be stable under ADTS. In contrast, if the data point is above the stability limit, it means that the beam will be unstable even if the ADTS is present.
Strictly speaking, the stability limit is the analytical solution to the dispersion relation [20, 21] where the growth rate of a bunched beam under ADTS is zero; . This stability boundary should agree with the tracking simulation results; see, e.g., [22].
@@
Figure 1 shows the stability diagram of the SOLEIL II booster at 150 MeV, where we can see the stability limit and transverse extracted from tracking simulations with different bunch charges. The horizontal axis is the mode tune shift where is the coherent tune obtained from the simulation, is the zero current coherent tune in the vertical plane, is the head-tail mode number, and is the synchrotron tune. This is a standard way to visualize the data, instead of showing the raw coherent tune shift, it shows the detuning from the corresponding head-tail mode in the unit of . In this plot, the charge is varied between . The data point for each chromaticity 111The chromaticities throughout this article are non-normalized chromaticities. evolves from the lower center of the diagram at a low charge, then climbs up vertically as the charge increases until it reaches the stability limit. It finishes in the upper left corner area at high charge at last. We can see that, at the nominal charge of 0.5 nC, the beam finds itself lying outside the stability limit with a growth rate in the order of while the maximum Landau damping rate is about .
@@
According to this result, the beam stability cannot be assumed for the rest of the ramp. However, as the head-tail growth rate decreases with energy, while the synchrotron radiation damping rate increases, it is possible that the instability will not appear at all. Therefore, simulations with the energy ramp are needed to evaluate the instability along the ramp.
@@ II.2 Along the ramp
We perform tracking simulations with mbtrack2, which uses the one-turn maps tracking method. A variation of beam energy is added to predict the realistic evolution of the beam throughout the ramp. The principle is to update, at each turn, the beam energy and equilibrium parameters such as energy loss per turn, radiation damping time, rf voltage and synchronous phase, natural energy spread, natural emittance, etc. Longitudinal short-range wakefields are always included in the simulations. Radiation damping time is included in all the simulations, unless we explicitly write that it has been excluded in one of the planes.The beam energy variation itself is calculated via the energy gain per turn equation in relation to the dipole field ramping rate, and the equilibrium parameters are calculated using their proportionality to the beam energy. More details on the booster ramp model can be consulted in [23, 10]. Moreover, the beam injected into the booster is not matched to the booster equilibrium parameters at the injection energy. Parameters of the injected beam, determined by the linac and the transfer line, are shown in Table 2.
@@
For the nominal charge in Fig. 2(a), it turns out that the beam is well stable and no sign of instability can be observed regardless of chromaticity. The vertical emittance is damped all the way from injection and reaches the designed value 0.5 nm rad at the extraction, which is the desirable situation. This result is rather surprising since some emittance increase could have been expected due to the high growth rate above the Landau stability limit shown in Fig. 1. The explanation for this is that even though Landau damping cannot totally suppress the instability, the growth rates are significantly reduced by the presence of ADTS, in addition to the reduction due to the increasing beam energy. This can be seen in Fig. 3 where growth rates extracted from the tracking simulations with and without ADTS are compared. For a certain chromaticity, the ADTS can greatly diminish the instability growth rate even at zero chromaticity where the Transverse Mode-Coupling Instability (TMCI) [24, 25, 26], which is extremely strong once emerged, takes place. At 0.5 nC for example, the growth rates can drop from to the minimum around which is less than or in the same order as the booster’s half-cycle rate . This means that the instability will never have enough time to build up before reaching the extraction, and the beam is simply damped by synchrotron radiation as the energy increases.
The growth rate reduction outside the stability limit can actually be viewed as a distortion of the complex tune shift space due to the presence of nonlinearity as we can see in Fig. 4. Here, the same stability limit shown previously is drawn in black. However, this is only a curve that corresponds to Im[]=0, the definition of a stability boundary. It means that the growth rate (under ADTS) along this curve is zero, and we can also call it an isoline of the zero growth rate. There are, in fact, an infinite number of isolines that show other growth rates Im[] (see e.g [27]). Here in Fig. 4, we plot some others showing Im[]=19.2, 96.6 and 191.2 . It can be clearly seen that the neighboring space of the stability limit is distorted. If we have a falling, for example, at the point marked as “x”, it means that the growth rate of without ADTS will be reduced to about under the ADTS.
@@ II.2.1 Sawtooth instability
Unique to the case of maximum charge, an unexpected sawtooth feature was observed at chromaticities 3 to 5. Further investigations were, therefore, carried out by simulating TSBI at different energies and the growth rates were compared to the analytical synchrotron radiation damping rate. It should be noted that, in this case, synchrotron radiation was not included in the transverse planes during the tracking in order to speed up the calculation and because we were interested in the pure instability growth rate. It was, however, included in the longitudinal plane to retain all the particles within the rf bucket222The rf bucket shrinks during the ramp, it is not an issue normally as the beam longitudinal emittance is also decreasing during the ramp due to radiation damping. But if the longitudinal synchrotron radiation damping is switched off, some particles will escape the rf bucket during the ramp, which disturbs the longitudinal binning profile used to compute longitudinal and transverse wake potentials..
The investigation results are shown in Fig. 5. Here, the growth rate of a 1.5 nC bunch at different chromaticities is determined at various points along the ramp discretely, meaning that each point represents an independent simulation with a fixed beam energy. We can see that the growth rates of the high chromaticity group (3 to 5) always remain above the radiation damping rate, while those of the low chromaticity group (0 to 2) drop below the latter at some point. Note that the growth rate at zero chromaticity has dropped below since the first 20 ms. As we have seen in Fig. 2(b) that only in the high chromaticity group did the sawtooth behavior occur, it is clear that the growth rate being stronger than radiation damping is the reason for the regrowing emittance. Another factor that supports this regrowth is the ineffectiveness of the ADTS at this point of the ramp due to the small emittance. However, as the emittance increases from the instability, the effect of ADTS becomes stronger and can damp the beam again. This mechanism repeats itself as long as the growth rate is larger than the radiation damping rate and it is the explanation to the sawtooth-like shape of the emittance.
Such an instability was also observed in the storage ring of APS in 1999 [28] and Landau damping was used to explain the self-limiting character [29]. More recently, MAX IV published a similar observation when the influence of ADTS on TMCI was being studied [30].
@@ II.2.2 Effect from the longitudinal damping
The results in Fig. 6 agree with the study in [32] where the effect of longitudinal damping on TCBI was investigated. In this reference, it is also shown that this effect is stronger when the longitudinal damping rate is larger. However, in our case, when we look at different chromaticities at the nominal longitudinal damping rate, the results do not follow the same conclusion as in this reference. We found that chromaticity 5 is the least affected by the longitudinal damping, and the lower the chromaticity, the stronger the effect. Meanwhile, in the reference, the effect of longitudinal damping is more important at higher chromaticity.
A benchmark on the growth rate as a function of chromaticity was done with a Vlasov solver DELPHI [33] and the result is shown in Fig. 7. Both codes were given with the same parameters and the same bunch length at each energy. The only difference is that DELPHI does not have radiation damping in any plane, while mbtrack2 still has one in the longitudinal plane. It evidently shows that the longitudinal radiation damping plays a role only up to chromaticity 2.
@@ III.1 General trend from analytical estimation
where is the average optic beta function, is the angular revolution frequency, is the total beam current, is the beam energy in eV, is the machine radius, is the chamber radius, is the speed of light, is the impedance of free space, is the fractional part of the betatron tune, and is the conductivity of the chamber wall. Apart from the vacuum chamber radius dependence, we can also see that the RWI growth rate is inversely proportional to the beam energy, . It should be mentioned that Eq. (3) was proved from a uniform beam filling pattern. Thus, in a case where the ring is partially filled, as ours, it can only give an estimated growth rate.
@@
is also shown. Note that, here, we are interested in the zero chromaticity circumstance, which can only provoke the mode . Theoretically, the longitudinal damping effect can then be neglected. A positive total growth rate means that instability is stronger than synchrotron radiation, and the beam center-of-mass (CM) oscillation can grow exponentially in the transverse plane, which is the sign of RWI. The point where the total growth rate goes into the negative side, therefore, marks the point where synchrotron radiation is strong enough to overtake RWI and possibly stabilize the beam. It can be seen in this figure that synchrotron radiation is ineffective during the first half of the ramp, resulting in possible instability.
To quickly predict the beam behavior, we define a simple equation to describe the beam CM transverse oscillation amplitude as a function of time :
@@
This result gives a general trend of RWI in every booster at zero chromaticity when there are no other stabilizing mechanisms involved, such as Landau damping, head-tail damping, or a transverse feedback system. The height of the curve depends on each machine’s ramp speed: the faster the ramp, the sooner synchrotron radiation can suppress RWI, and the smaller the peak oscillation will be. Another important factor is the vacuum chamber radius. For 3GLS boosters whose chamber radius is typically large, the RWI growth rate is small, so the CM oscillation is relatively small as a consequence.